Could you please tell me the answer to this question: What is the equation of a line that passes through the points (0,3) and (-2,-3)? I know the equation for a line is (y=mx + b) but I'm unsure on how to do it. Thank you very much :D
Answer by Theo(13342)

the line passes through the points (0,3) and (-2,-3)

those points are in (x,y) format.

the slope of a straight line is (y2-y1)/(x2-x1)

assign one of the points to (x1,y1) and assign the other point to (x2,y2)

i chose (x1,y1) = (0,3) and (x2,y2) = (-2,-3).

you could have chosen it the other way around ((x1,y1) = (-2,-3) and (x2,y2) = (0,3)).

you'll get the same slope either way.

the slope intercept form of the line is y = mx + b

m is the slope
b is the y-intercept.

the formula for the slope, given 2 points is (y2-y1)/(x2-x1)

replace (x1,y1) and (x2,y2) with their respective values, and the formula becomes:

when = (x1,y1) = (0,3) and (x2,y2) = (-2,-3), the slope formula becomes:

slope = (-3-3) / (-2-0) which becomes -6/-2 which becomes 3.

the slope is equal to 3.

since m is the slope, replace m with 3 in the equation of y = mx + b to get y = 3x + b

now all that is left is to find the y-intercept.

replace x and y with the value of one of the points.

either one will do.

i chose (x2,y2) = (-2,-3)

y = 3x + b becomes -3 = 3*-2 + b
this becomes -3 = -6 + b
add 6 to both sides of this equation to get 3 = b

the equation of y = 3x + b becomes y = 3x + 3

the slope is 3 and the y-intercept is 3.

here's a graph of the equation.

you can see that the point (-2,-3) is on the line and the point (0,3) is also on the line.

$$$

here's a reference you might find helpful.

http://www.purplemath.com/modules/strtlneq.htm

some tutorials tell you to use the point slope form of the equation to find the y-intercept.

since we already found the slope, our equation is y = 3x + b

to find the y-intercept, use the slope intercept form of the equation of a straight line.

that form is y-y1 = m * (x-x1)

x1 and y1 are any one of the two points you were given.

we'll use (0,3)

the equation becomes y - 3 = 3 * (x-0)

simplify to get y - 3 = 3x

add 3 to both sides to get y = 3x + 3

3 is the y-intercept.

a reference on point slope form of the equation of a straight line can be found here.

http://www.purplemath.com/modules/strtlneq2.htm